What is the pH of 44.3 ml of a solution which is 0.26 M in NaF and 0.33 M in HF?

For HF use Ka=6.8x10^-4
Answer #1

Ka = 6.8*10^-4

pKa = - log (Ka)

= - log(6.8*10^-4)

= 3.167

use:

pH = pKa + log {[conjugate base]/[acid]}

= 3.167+ log {0.26/0.33}

= 3.064

Answer: 3.06